OpenStudy (anonymous):

Write the slope intercept equation for the line with slope -1/4 and y intercept 0,-2

OpenStudy (angela210793):

y=mx+k y=-x/4-2

OpenStudy (anonymous):

use y-y1=m(x-x1) where m= slope